# Compact Minimal Submanifolds of the Riemannian Symmetric Spaces

### v1

@@ 11. The Symmetric Space $\text{\bf SU}(2n)/\text{\bf Sp}(n)$
In this section we prove Theorem 11.2 and thereby construct a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\text{\bf SU}(2n)/\text{\bf Sp}(n)$. It is well-known that the natural projecton $\pi:\text{\bf SU}(2n)\to\text{\bf SU}(2n)/\text{\bf Sp}(n)$ from the special unitary group $\text{\bf SU}(2n)$ is a Riemannian submersion. This means that we can employ Corollary 2.6 in this situation.
@@
We now state Theorem 11.2, which is our main result of this section.
Theorem 11.2.
$A=\sum_{j,\alpha=1}^{2n}a_{j}b_{\alpha}Y_{j\alpha}.$
$\hat{\phi}_{a,b}(z)=\operatorname{trace}(AzJ_{n}z^{t}).$
@@
Let us assume that $z\in\text{\bf SU}(2n)$ is an element of the fibre $\hat{\phi}_{a,b}^{-1}(\{0\})$ i.e. $\hat{\phi}_{a,b}(z)=0$. The statement of Corollary 3.2 guarantees the existence of such a point $z$. Further assume that $z$ is a critical point of $\hat{\phi}_{a,b}:\text{\bf SU}(2n)\to{\mathbb{C}}$ which means that $Y(\hat{\phi}_{a,b}(z))=0$ for every element in the orthogonal complement $\mathfrak{p}$ of $\mathfrak{sp}(n)$ in $\mathfrak{su}(2n)$. It can now be shown that for all $1\leq r<s\leq n,$
$Y_{rs}^{a}(\phi_{a,b})=0,\ Y_{rs}^{b}(\phi_{a,b})=0,\ X_{rs}^{a}(\phi_{a,b})=0,\ Y_{rs}^{c}(\phi_{a,b})=0$
@@
Without loss of generality, we can assume that $1\leq t\leq n.$ Equations 11.1 and 11.3 now show that for all $1\leq s\leq n,\ s\neq t,$ it holds that
@@ 12. Acknowledgements
The authors are grateful to Thomas Jack Munn for useful discussions on this work.
@@ Appendix A The linear system for $\mathbf{SO}(2n)/\text{\bf U}(n)$
The matrix $M_{2}$ can then be transfomed into $\tilde{M}_{2}=(u,x_{1})\cdot I_{4}+S_{2}$.
@@
The matrix $M_{3}$ can then be transfomed into $\tilde{M}_{3}=(u,x_{1})\cdot I_{6}+S_{3}$.
@@
The corresponding matrix $M_{n}$ can then be transfomed into $\tilde{M}_{n}=(u,x_{1})\cdot I_{2n}+S_{n}$. Here the matrix $M_{n}\in{\mathbb{R}}^{2n\times 2n}$ satisfies
$|\det M_{n}|=(\,(u,x_{1})^{2}+(u,x_{n+1})^{2}\,)^{n-2}\cdot|u|^{4}\neq 0.$

### v2

@@ 12. The Symmetric Space $\text{\bf SU}(2n)/\text{\bf Sp}(n)$
In this last section we provide the proof of Theorem 2.4. Consequently, we obtain a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\text{\bf SU}(2n)/\text{\bf Sp}(n)$. Similar to the previous sections, note that the natural projection $\pi:\text{\bf SU}(2n)\to\text{\bf SU}(2n)/\text{\bf Sp}(n)$ from the special unitary group $\text{\bf SU}(2n)$ is a Riemannian submersion and thus Corollary 3.5 applies.
@@
(Theorem 2.4) Let us assume that $z\in\text{\bf SU}(2n)$ is an element of the fibre $\hat{\phi}_{a,b}^{-1}(\{0\})$ i.e. $\hat{\phi}_{a,b}(z)=0$. The statement of Corollary 4.3 guarantees the existence of such a point $z$. Further assume that $z$ is a critical point of $\hat{\phi}_{a,b}:\text{\bf SU}(2n)\to{\mathbb{C}}$ which means that $Y(\hat{\phi}_{a,b}(z))=0$ for every element in the orthogonal complement $\mathfrak{p}$ of $\mathfrak{sp}(n)$ in $\mathfrak{su}(2n)$. It can now be shown that for all $1\leq r<s\leq n,$
$Y_{rs}^{a}(\hat{\phi}_{a,b})=0,\ Y_{rs}^{b}(\hat{\phi}_{a,b})=0,\ X_{rs}^{a}(\hat{\phi}_{a,b})=0,\ Y_{rs}^{c}(\hat{\phi}_{a,b})=0$
@@
Without loss of generality, we can assume that $1\leq t\leq n.$ Equations (12.1) and 12.3 now show that for all $1\leq s\leq n,\ s\neq t,$ it holds that
@@ 13. Acknowledgements
The authors are grateful to Thomas Jack Munn and Oskar Riedler for useful discussions on this work. They would also like to thank the anonymous referee for important suggestions on the presentation.
@@ Appendix A The linear system for $\mathbf{SO}(2n)/\text{\bf U}(n)$
The matrix $M_{2}$ can then be transformed into $\tilde{M}_{2}=(u,x_{1})\cdot I_{4}+S_{2}$.
@@
The matrix $M_{3}$ can then be transformed into $\tilde{M}_{3}=(u,x_{1})\cdot I_{6}+S_{3}$.
@@
The corresponding matrix $M_{n}$ can then be transformed into
$\tilde{M}_{n}=(u,x_{1})\cdot I_{2n}+S_{n}.$